OpenStudy (anonymous):

3. Factor the following trinomials: x2 + 10x + 16

myininaya (myininaya):

\[(x+8)(x+2)\]

OpenStudy (mimi_x3):

(x+2)(x+8)

myininaya (myininaya):

since 8(2)=16 and 8+2=10

OpenStudy (anonymous):

(x+2)(x+8)

OpenStudy (anonymous):

\[x^2+10x+16\] \[x^2+8x+2x+16\] \[x(x+8) + 2(x+8)\] this equation is now same as previous one a(b+c) + d(b+c) = (a+d)(b+c) therefore \[(x+2)(x+8)\]